# Gaussian Processes enabled model calibration in the context of deep geological disposal

### v1

@@ Abstract
This work introduces a surrogate modeling approach for an emplacement drift of a deep geological repository based on Gaussian Processes (GPs). The surrogate model is used as a substitute for the high-fidelity mechanical model in many-query scenarios, such as time-dependent sensitivity analysis and calibration. Our GP-based approach emulates the behavior of an emplacement drift of a deep geological repository with significantly reduced computational time, enabling faster design iterations and effective incorporation as well as interpretation of monitoring data. Our findings show that only a few key parameters are essential for accurately reflecting in-situ conditions in complex rock salt models, which is critical for ensuring safety in deep geological disposal.
@@
articletype: Research article
@@ 1 Introduction
Deep repository models are complex geological models that account for the mechanics of soil and rock, hydrological properties, thermal effects, and chemical interactions. They are highly parametrized and the associated numerical analyses are computationally expensive. Their complexity has prevented their widespread adoption for many-query tasks, that is, simulations to explore multiple scenarios and the effects of uncertainties on the deep repository prognosis. In this regard, we propose using Gaussian Processes (GPs) as surrogates for high-fidelity geological models. GPs can approximate the outputs of complex simulations with much lower computational cost, enabling efficient calibration, design and validation. Speeding up the simulations with the use of a GP-based surrogate will not only allow for validated prognoses for the repository through the assimilation of data but also help in making informed decisions and enhancing the overall robustness and adaptability of the repository management process.
@@ 1.2 Gaussian processes as efficient surrogates for computationally expensive models
In [40] the concept of using GPs as surrogates to predict model outputs at untried locations within the parameter space was introduced. The use of GPs for calibrating computer models was then pioneered by [27] and [28]. Subsequent work by [6] and [22] refined this framework, with Higdon adopting a fully Bayesian approach. Extending these methodologies to multivariate data [23], and further advancements by others, such as [5], have enhanced the calibration of models with multivariate outputs. [43], and additional studies, including those by [11], [32], and [4], provide in-depth discussions on the optimal design of computer experiments, focusing on the strategic layout of model evaluations in parameter space. Beyond calibration, GP emulators facilitate understanding variability in model outputs when parameters are uncertain, known as uncertainty analysis. The tutorial by [38] offers an accessible introduction to uncertainty analysis using GP emulators.
1.3 Paper contributions and outline
The primary contributions of this paper are:
@@ 2.1 Detailed site description and problem geometry
Figure 1: Vertical geological cross-section of the Gorleben salt dome according to [7], showing the location of the considered drift (small red circle).
The considered measurement cross-section is located in the northern main drift of Gorleben with a depth of $840$ meters below the top edge of the ground. Excavation in the area of the measurement location was finished on October 19, 1999 without a recut being carried out afterwards. The monitoring data consists of time series of convergence measurements, which indicate the change in distance between opposing fixed points inside the rock, from which the deformation rate of the drift contour is derived. At each measuring location, the horizontal and vertical distances are recorded periodically as a standard procedure. Consequently, the computational model adopted in this work represents a similar open emplacement drift of a deep geological repository based on the location where the monitoring data were obtained. Since this phase does not involve storing highly radioactive and heat-generating waste, the problem considered is purely mechanical.
@@
Figure 2 (a) depicts the cross-section of the measurement location together with the vertical (2-4) and horizontal (1-3) measurement distances. Figure 2 (b) corresponds to the computational model of the drift in FLAC3D with history locations for the evaluation of the displacements.
Figure 2: Cross-sectional representation of the drift area. (a) Cross-section of the measurement locations from the Gorleben site. (b) Computational model of the drift in FLAC3D, showing the history locations used for evaluating displacements.
@@ 2.2 Continuum mechanics
Modeling in continuum mechanics involves three ingredients: i. balance equations, ii. kinematics and iii. a constitutive model. This section introduces these three ingredients of continuum mechanics for the deep repository under the assumptions presented in the previous Section 2.1.
@@ 2.2.1 Balance equations and kinematics
The boundary value problem is complemented by the constitutive model TUBSsalt, to be introduced in Section 2.2.2, and solved using the commercial software FLAC3D [25] as will be discussed in Section 2.3. Only small deformations are considered in this work since no long-term analysis of the rock salt emplacement will be performed. Therefore, the kinematics are defined by
@@ 2.2.2 Constitutive model TUBSsalt
Figure 3: Rheological model of the constitutive model TUBSsalt for rock salt according to [18].
@@
Here, $\dot{\bm{\varepsilon}}_{el}$ is the elastic strain rate and $\dot{\bm{\varepsilon}}_{vp}$ the inelastic, visco-plastic strain rate. The latter consists of the primary creep rate $\dot{\bm{\varepsilon}}_{p}$, the secondary creep rate $\dot{\bm{\varepsilon}}_{s}$, the tertiary creep and healing rate $\dot{\bm{\varepsilon}}_{t}$, the strain rate from creep and shear failure $\dot{\bm{\varepsilon}}_{n}$ and the strain rate from tension failure $\dot{\bm{\varepsilon}}_{z}$. In Figure 3, the elastic strain rate $\dot{\bm{\varepsilon}}_{el}$ is represented by a spring, and the creep strain rates $\dot{\bm{\varepsilon}}_{p}$, $\dot{\bm{\varepsilon}}_{s}$ and $\dot{\bm{\varepsilon}}_{t}$ as well as the failure strain rates $\dot{\bm{\varepsilon}}_{n}$ and $\dot{\bm{\varepsilon}}_{z}$ are described by hardening or softening sliders with yield functions and viscous dampers. For a visco-elasto-plastic material model, the stress increment depends on the elastic strain component and can be written as
@@
For further details on quantities not explained in this section, such as the yield functions $F_{t}$, $F_{p}$, $F_{n}$, and the potential functions for tertiary creep and healing $Q$ and $Q_{v}$, the reader is referred to [10]. In summary, the constitutive model TUBSsalt comprises a total of 25 material parameters, summarized in Appendix B, Table 1.
@@ 2.3 Numerical solution
The constitutive model TUBSsalt introduced in Section 2.2.2 is implemented into the commercial software FLAC3D [25], which is a program for three-dimensional engineering mechanics computations. FLAC3D relies on the finite volume method to translate the continuum equations into ordinary differential equations for each element, which are then solved using an explicit central finite difference approach in time. The spatial discretization in FLAC3D is realized by meshing the continuum into hexahedral elements, with the option to use tetrahedral, wedge, and pyramid elements. It is recommended to use hexahedral elements, which are divided into tetrahedral elements to reduce the volumetric locking effect and thus achieve more accurate results. The balance of the linear momentum (2.1) presented in Section 2.2 is iterated to an equilibrium state, which is achieved when the unbalanced mechanical force for all the gridpoints in the model negligibly low.
Time dependent phenomena such as creep require a timestep $\Delta t$ to solve the equations of the constitutive model. At the same time, for creep analysis, the state of equilibrium must be maintained, otherwise inertial effects may affect the solution. For this purpose, the unbalanced force is monitored in the model. The finite volume scheme can be summarised for each time step as follows: After new strain rates are determined from nodal velocities, new stresses are calculated from strain rates and previous stresses using constitutive equations. By applying the balance of the linear momentum, new velocities and deformations are subsequently calculated from stresses and forces. More information of the FLAC3D solution algorithm can be found in [25].
@@ 3.1 Initial model calibration from mechanical testing data
Commonly, laboratory tests are used for the calibration of constitutive model parameters. In the case of rock salt and as shown in Table 1, short-term triaxial compression or extension tests, long-term creep tests and healing tests as well as indirect tensile tests need to be conducted under variation of temperature and confining stress to be able to cover all strain parts described in Section 2.2.2. While strength tests use a constant axial strain rate to apply a load on the mostly cylindrical specimens with a fixed confining pressure, the specimens in creep and healing tests have a predefined stress state so that the time-dependent behavior can be observed.
@@
However, the samples used in laboratory tests are no longer in their original state and the material tests only give insight into the material behavior in localized regions from where the sample has been extracted. As a consequence, model predictions typically do not match with real-world monitoring data, necessitating re-calibration of the constitutive model. Therefore, model calibration based on monitoring data is addressed in the following Section 3.2.
@@ 3.2 Inverse problem formulation for the model re-calibration
Let $\mathbf{Y}_{\text{moni}}^{T}=[\mathbf{y}_{\text{moni}}^{T}(t_{1}),\,\mathbf{y}_{\text{moni}}^{T}(t_{2}),\,...,\,\mathbf{y}_{\text{moni}}^{T}(t_{n})]$ denote the vector of experimental observations at different time instances for a specific location with $n$ the total number of time instances. The corresponding model prediction is denoted by $\mathbf{s}\left(t_{i};\bm{\theta}\right)$. Then, in analogy to equation (3.1), material parameter are identified from
$\bm{\theta}^{*}=\text{arg min}_{\bm{\theta}}\ \sum_{i=1}^{n}\|\mathbf{y}_{\text{moni}}(t_{i})-\mathbf{s}(t_{i},\bm{\theta})\|^{2}.$ (3.2)
Note that in (3.1) and (3.2) we omit a weighting matrix, that is often introduced to account for the size of measurement errors. In our setting, the error sizes did not have a relevant influence. The optimization problem (3.2) is implemented using the SciPy Python library [57], employing a global optimization algorithm. We opted for the differential evolution algorithm [50], but also experimented with genetic and dual annealing algorithms. However, differential evolution proved to be faster and more reliable in this context.
@@ 3.3 Sensitivity analysis based on time-dependent Sobol’ indices
In this contribution, we compute sensitivities based on Sobol’ indices [46, 41, 42]. The first-order Sobol’ index quantifies the contribution of each parameter $\theta_{j}$ to the variance in the predicted solution $\mathbf{s}_{i}(\bm{\theta})=\mathbf{s}(t_{i},\,\bm{\theta})$, while averaging out the effects of other inputs. The first-order Sobol’ index for each time point $t_{i}$ is defined as:
@@ 4 Gaussian processes as surrogate models for complex computer models
The model $\mathbf{s}$ represents the mechanical behavior of the deep geological repository introduced in Section 2. Evaluating $\mathbf{s}$ directly is computationally demanding, especially in many-query scenarios such as model calibration, optimization, or design. Surrogate modeling offers an efficient alternative to address the computational challenges associated with directly solving $\mathbf{s}$. Among the available techniques, Proper Orthogonal Decomposition (POD) [1] and Physics-Informed Neural Networks (PINNs) [3] are particularly suitable for full-field simulations where capturing the entire spatial domain is crucial. However, GPs present a compelling option when the focus is on specific quantities of interest, such as an excavation convergence as in the present study, rather than the entire field. GPs are especially advantageous due to their ability to provide probabilistic predictions and handle smaller datasets effectively. In this study, we adopt GPs to develop a surrogate model, denoted as $\hat{\mathbf{s}}$, to approximate the complex behavior of $\mathbf{s}$.
@@ 5.1 High-fidelity model simulation
Monitoring data: The mechanical model described in Section 2 represents an open drift located in the northern main drift of Gorleben for which monitoring data is available. The first convergence measurement took place on October 20, 1999, one day after the excavation of the drift. This is important because it allows for the measurement of primary creep, which is most pronounced at the beginning of the excavation. On the day of the initial measurement, the horizontal distance (1-3) measured $8.94$ meters, and the vertical distance (2-4) was $6.00$ meters, as depicted in Figure 2(a). The duration of the monitoring was approximately $14.3$ years, with a significant variation of time intervals between measurements. During the first two months, measurements were taken every two to three days in order to capture the high creep rates. Subsequently, the measurement interval was increased from one week to one month, and from the end of 2002 onwards, measurements were taken approximately every six months as a constant creep rate had been achieved. Since the measuring points of the convergence distances are anchored in the rock over an anchoring length, the deformations of the cavity are also measured at these fixed points in the rock salt. Those locations are defined as history points in the numerical model in FLAC3D, demonstrated in Figure 2(b). The convergences $\mathbf{y}_{\text{moni}}^{T}(t_{i})=\left[\Delta u_{x}(t_{i}),\,\Delta u_{z}(t_{i})\right]^{T}$ are measured as a vertical and horizontal distance between two points, therefore the displacements $u_{x}$ in horizontal and $u_{z}$ in vertical direction are obtained from the displacement vector $\mathbf{u}$ for the specific nodes.
@@
Preliminary parameter selection: TUBSalt has a total of $25$ material parameters. Out of these, $5$ parameters can be read directly from laboratory stress-strain data, namely $K_{0}$, $G_{0}$, $\varepsilon_{v,d,b}$, $\sigma_{z,0}$ and $\rho$. In the following, these are considered as well-calibrated and independent of the monitoring data. The remaining $20$ parameters serve as potential parameters for calibration. We then conducted a preliminary, empirical study with only two simulations: several damage-associated strain components were deactivated, and the result was compared to a reference solution obtained with the full set of strain components (2.5). It has been found that the test case considered is dominated by primary and secondary creep mechanisms and damage-associated strain components are negligible. Details can be found in Appendix A. As a result, in the following we focus on a total of $7$ TUBSsalt parameters for primary and secondary creep: $\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$ and $E_{p}$ for primary creep and $\eta_{s}$, $\sigma_{0,eq,s}$ and $p_{s}$ for secondary creep, and thus $\bm{\theta}=\left[\eta_{p},\,\sigma_{0,eq,p},\,p_{p},\,E_{p},\,\eta_{s},\,\sigma_{0,eq,s},\,p_{s}\right]$. The parameter ranges of those $7$ parameters, as well as the parameters that are considered as fixed, are summarized in Appendix B, Table 1.
@@
Sampling of parameter values in Flac3D: Following the approach in Section 3.1, a parameter set for the constitutive model TUBSsalt based on laboratory strength and creep tests has been determined in a project presented in [48] for the Gorleben site. These values serve as characteristic parameter values for Gorleben salt and also help define reasonable parameter ranges for those parameters that need calibration, see also Table 1 in Appendix B. In order to create training and testing data by means of a multitude of simulations, the parameter values of the selected TUBSsalt parameters are sampled in the given ranges in Table 1 assuming a uniform distribution. For this purpose, a SciPy Quasi-Monte Carlo generator is used to generate a Sobol sequence by creating low-discrepancy, quasi-random numbers [45, 37]. The loop for parameter sampling is implemented in the creep analysis of Flac3D, so that the saved equilibrium state of the model is recalled for every iteration. The creep analysis is performed for the same duration of the monitoring. To maintain the state of equilibrium, the average ratio between out-of-balance and total forces is limited to $5\cdot 10^{-6}$. The timestep increases over time with a maximum of $\Delta t=5.56$ hours. With the given configuration, a single simulation run of the high-fidelity model requires approximately $4$ minutes.
@@
Processing the full-field simulation results to convergences: After carrying out the simulations to create training and testing data, the displacements derived at the history locations are converted into convergences at the time instances for which monitoring data are available. To determine the vertical convergences, the amounts of the vertical displacements $u_{z}(t_{i})$ at the top and bottom are summed up together. To determine the horizontal convergences, the amount of the horizontal displacements $u_{x}(t_{i})$ is multiplied by two to account for the symmetry of the system. The displacements of an exemplary full-field simulation at $t=14.5$ years is depicted in Figure 4. In order to compare the simulated convergences with the monitoring data, the convergences of the simulations must be set to zero at the time of the initial monitoring measurement. Therefore, the convergences from the first day after excavation of the drift can not be analysed and will be neglected. Convergences are determined at all points in time at which monitoring is available. As a result Figure 5 shows the processed data of $200$ simulations (blue lines) over time in comparison with the corresponding monitoring data (black squares) for both horizontal and vertical convergences.
(a) Vertical displacements $\mathbf{u}_{z}$
(b) Horizontal displacements $\mathbf{u}_{x}$
@@
Figure 5: Simulation over time at the monitoring location of the vertical (left) and horizontal (right) convergences by varying parameters $\eta_{p}$, $E_{p}$, $\eta_{s}$, $p_{s}$, $\sigma_{0,eq,p}$, $p_{p}$, and $\sigma_{0,eq,s}$ of TUBSsalt. The plot shows monitoring data (black squares) and $200$ model realizations (blue lines) from the FLAC3D simulations.

### v2

@@ Abstract
Deep geological repositories are critical for the long-term storage of hazardous materials, where understanding the mechanical behavior of emplacement drifts is essential for safety assurance. This study presents a surrogate modeling approach for the mechanical response of emplacement drifts in rock salt formations, utilizing Gaussian Processes (GPs). The surrogate model serves as an efficient substitute for high-fidelity mechanical simulations in many-query scenarios, including time-dependent sensitivity analyses and calibration tasks. By significantly reducing computational demands, this approach facilitates faster design iterations and enhances the interpretation of monitoring data. The findings indicate that only a few key parameters are sufficient to accurately reflect in-situ conditions in complex rock salt models. Identifying these parameters is crucial for ensuring the reliability and safety of deep geological disposal systems.
@@
journal: Data/Matharticletype: Research article
[…]
2024
@@ 1 Introduction
Deep repository material models are complex geological models that account for the mechanics of soil and rock, hydrological properties, thermal effects, and chemical interactions [1, 2, 3]. They are highly parametrized and the associated numerical analyses are computationally expensive [4, 5]. Their complexity has prevented their widespread adoption for many-query tasks, that is, simulations to explore multiple scenarios and the effects of uncertainties on the deep repository prognosis [5]. In this regard, we propose using Gaussian Processes (GPs) as surrogates for high-fidelity geological models. GPs can approximate the outputs of complex simulations with much lower computational cost, enabling efficient calibration, design and validation [6, 7, 8]. Speeding up the simulations with the use of a GP-based surrogate will not only allow for validated prognoses for the repository through the assimilation of data but also help in making informed decisions and enhancing the overall robustness and adaptability of the repository management process.
@@ 1.2 Gaussian processes as efficient surrogates for computationally expensive models
In [35], the concept of using GPs as surrogates to predict model outputs at untried locations within the parameter space was introduced. The use of GPs for calibrating computer models was then pioneered by [17] and [18]. Subsequent work by [36] and [37] refined this framework, with Higdon adopting a fully Bayesian approach. Extending these methodologies to multivariate data [38], and further advancements by others, such as [39], have enhanced the calibration of models with multivariate outputs. [40], and additional studies, including those by [41], [42], and [43], provide in-depth discussions on the optimal design of computer experiments, focusing on the strategic layout of model evaluations in parameter space. Beyond calibration, GP emulators facilitate understanding variability in model outputs when parameters are uncertain, known as uncertainty analysis. The tutorial by [44] offers an accessible introduction to uncertainty analysis using GP emulators.
Unlike much of the existing literature on surrogate modeling for complex mechanical systems, which often relies on synthetic data, our study goes beyond the implementation of a surrogate modeling approach and exploits it to enable the calibration of the TUBSsalt constitutive model using 14 years of real-world monitoring data collected in an open drift located in the northern main drift of Gorleben, Germany. The primary contribution of this paper is to demonstrate the applicability and effectiveness of a GP-based data-driven methodology in addressing a real-world problem. Specifically:
@@ 2.1 Detailed site description and problem geometry
Figure 1: Vertical geological cross-section of the Gorleben salt dome, showing the location of the considered drift (small red circle). The stratigraphy was obtained from the open-access report [45].
The considered measurement cross-section is located in the northern main drift of Gorleben with a depth of $840\,m$ below the top edge of the ground. Excavation in the area of the measurement location was finished on October 19, 1999 without a recut being carried out afterwards. The monitoring data consists of time series of convergence measurements, which indicate the change in distance between opposing fixed points inside the rock, from which the deformation rate of the drift contour is derived. At each measuring location, the horizontal and vertical distances are recorded periodically as a standard procedure. Consequently, the computational model adopted in this work represents a similar open emplacement drift of a deep geological repository based on the location where the monitoring data were obtained. Since this phase does not involve storing highly radioactive and heat-generating waste, the problem considered is purely mechanical.
@@
Figure 2a depicts the cross-section of the measurement location together with the vertical (2-4) and horizontal (1-3) measurement distances. Figure 2b corresponds to the computational model of the drift in FLAC3D with history locations for the evaluation of the displacements.
Figure 2: Cross-sectional representation of the drift area. a. Cross-section of the measurement locations from the Gorleben site, provided by the BGE mbH. b. Computational model of the drift in FLAC3D, showing the history locations used for evaluating displacements.
@@ 2.2 Continuum mechanics
Modeling in continuum mechanics involves three ingredients: i. balance equations, ii. kinematics and iii. a constitutive model. This section introduces these three ingredients of continuum mechanics for the deep repository under the assumptions presented in the previous Section. We follow the notation for general continuum mechanical equations introduced by [47, 48]. The constitutive equations for TUBSsalt are presented as they were introduced in [16, 49].
@@ 2.2.1 Balance equations and kinematics
The boundary value problem is complemented by the constitutive model TUBSsalt and will be solved using the commercial software FLAC3D [48].
Only small deformations are considered in this work since no long-term analysis of the rock salt emplacement will be performed. Therefore, the kinematics are defined by
@@ 2.2.2 Constitutive model TUBSsalt
Figure 3: Rheological model and corresponding strain components of the constitutive model TUBSsalt for rock salt. While $\dot{\bm{\varepsilon}}_{el}$ is represented by a spring, $\dot{\bm{\varepsilon}}_{p}$, $\dot{\bm{\varepsilon}}_{s}$, $\dot{\bm{\varepsilon}}_{t}$, $\dot{\bm{\varepsilon}}_{n}$ as well as $\dot{\bm{\varepsilon}}_{z}$ are modelled by hardening or softening sliders and viscous dampers.
@@
Here, $\dot{\bm{\varepsilon}}_{el}$ is the elastic strain rate and $\dot{\bm{\varepsilon}}_{vp}$ the inelastic, visco-plastic strain rate. The latter consists of the primary creep rate $\dot{\bm{\varepsilon}}_{p}$, the secondary creep rate $\dot{\bm{\varepsilon}}_{s}$, the tertiary creep and healing rate $\dot{\bm{\varepsilon}}_{t}$, the strain rate from creep and shear failure $\dot{\bm{\varepsilon}}_{n}$ and the strain rate from tension failure $\dot{\bm{\varepsilon}}_{z}$. In Figure 3, the elastic strain rate $\dot{\bm{\varepsilon}}_{el}$ is represented by a spring, and the creep strain rates $\dot{\bm{\varepsilon}}_{p}$, $\dot{\bm{\varepsilon}}_{s}$ and $\dot{\bm{\varepsilon}}_{t}$ as well as the failure strain rates $\dot{\bm{\varepsilon}}_{n}$ and $\dot{\bm{\varepsilon}}_{z}$ are described by hardening or softening sliders with yield functions $F$ and viscous dampers, each characterised by an individual viscosity $\eta$. For a visco-elasto-plastic material model, the stress increment depends on the elastic strain component and can be written as
@@
Tension failure occurs when tensile material strength $\sigma_{z,0}$ is exceeded and is denoted by the strain rate tensor $\dot{\bm{\varepsilon}}_{z}$.
@@
For further details on quantities not explained in this section, such as the yield functions $F_{p}$, $F_{t}$, $F_{n}$, the potential functions for tertiary creep and healing $Q$ and $Q_{v}$, as well as derivations of all equations of TUBSsalt provided here, the reader is referred to [49]. In summary, the constitutive model TUBSsalt comprises a total of 25 material parameters, summarized in Appendix B, Table B1.
@@ 2.3 Numerical solution
The constitutive model TUBSsalt is implemented into the commercial software FLAC3D [48], which is a program for three-dimensional engineering mechanics computations. FLAC3D relies on the finite volume method to translate the continuum equations into ordinary differential equations for each element, which are then solved using an explicit central finite difference approach in time. The spatial discretization in FLAC3D is realized by meshing the continuum into hexahedral elements, with the option to use tetrahedral, wedge, and pyramid elements. It is recommended to use hexahedral elements, which are divided into tetrahedral elements to reduce the volumetric locking effect and thus achieve more accurate results. The balance of the linear momentum, equation (2.1), is iterated to an equilibrium state, which is achieved when the unbalanced mechanical force for all the gridpoints in the model is negligibly low.
Time-dependent phenomena such as creep require a timestep $\Delta t$ to solve the equations of the TUBSsalt constitutive model. At the same time, for creep analysis, the state of equilibrium must be maintained, otherwise inertial effects may affect the solution. For this purpose, the unbalanced force is monitored in the model. The finite volume scheme can be summarised for each time step as follows: After new strain rates are determined from nodal velocities, new stresses are calculated from strain rates and previous stresses using constitutive equations. By applying the balance of the linear momentum, new velocities and deformations are subsequently calculated from stresses and forces. More information on the FLAC3D solution algorithm can be found in [48].
@@ 3.1 Initial model calibration from mechanical testing data
Commonly, laboratory tests are used for the calibration of constitutive model parameters. In the case of rock salt and as shown in Table B1, short-term triaxial compression or extension tests, long-term creep tests and healing tests as well as indirect tensile tests need to be conducted under variation of temperature and confining stress to be able to cover all strain parts described in Section 2.2.2. While strength tests use a constant axial strain rate to apply a load on the mostly cylindrical specimens with a fixed confining pressure, the specimens in creep and healing tests have a predefined stress state so that the time-dependent behavior can be observed.
@@
However, the samples used in laboratory tests are no longer in their original state and the material tests only give insight into the material behavior in localized regions from where the sample has been extracted. As a consequence, model predictions typically do not match with real-world monitoring data, necessitating re-calibration of the constitutive model. Therefore, model calibration based on monitoring data is addressed in the following Section.
@@ 3.2 Inverse problem formulation for the model re-calibration
Let $\mathbf{Y}_{\text{moni}}^{T}=[\mathbf{y}_{\text{moni}}^{T}(t_{1}),\,\mathbf{y}_{\text{moni}}^{T}(t_{2}),\,...,\,\mathbf{y}_{\text{moni}}^{T}(t_{n})]$ denote the vector of experimental observations at different time instances for a specific location with $n$ the total number of time instances. The corresponding model prediction is denoted by $\mathbf{s}_{i}(\bm{\theta})=\mathbf{s}(t_{i},\,\bm{\theta})$. Then, in analogy to equation (3.1), material parameter are identified from
$\bm{\theta}^{*}=\text{arg min}_{\bm{\theta}}\ \sum_{i=1}^{n}\|\mathbf{y}_{\text{moni}}(t_{i})-\mathbf{s}_{i}(\bm{\theta})\|^{2}.$ (3.2)
Note that in equations (3.1) and (3.2) we omit a weighting matrix, that is often introduced to account for the size of measurement errors. In our setting, the error sizes did not have a relevant influence. The optimization problem (3.2) is implemented using the SciPy Python library [51], employing a global optimization algorithm. We opted for the differential evolution algorithm [52], but also experimented with genetic and dual annealing algorithms. However, differential evolution proved to be faster and more reliable in this context.
@@ 3.3 Sensitivity analysis based on time-dependent Sobol’ indices
In this contribution, we compute sensitivities based on Sobol’ indices [53, 54, 55]. The first-order Sobol’ index quantifies the contribution of each parameter $\theta_{j}$ to the variance in the predicted solution $\mathbf{s}_{i}(\bm{\theta})$, while averaging out the effects of other inputs. The first-order Sobol’ index for each time point $t_{i}$ is defined as:
@@ 4 Gaussian processes as surrogate models for complex computer models
The model $\mathbf{s}_{i}(\bm{\theta})$ represents the mechanical behavior of the deep geological repository introduced in Section 2. Evaluating $\mathbf{s}_{i}(\bm{\theta})$ directly is computationally demanding, especially in many-query scenarios such as model calibration, optimization, or design. Surrogate modeling offers an efficient alternative to address the computational challenges associated with directly solving $\mathbf{s}_{i}(\bm{\theta})$. Among the available techniques, Proper Orthogonal Decomposition (POD) [60] and Physics-Informed Neural Networks (PINNs) [61] are particularly suitable for full-field simulations where capturing the entire spatial domain is crucial. However, GPs present a compelling option when the focus is on specific quantities of interest, such as an excavation convergence as in the present study, rather than the entire field. GPs are especially advantageous due to their ability to provide probabilistic predictions and handle smaller datasets effectively. In this study, we adopt GPs to develop a surrogate model, denoted as $\hat{\mathbf{s}}_{i}(\bm{\theta})$, to approximate the complex behavior of $\mathbf{s}_{i}(\bm{\theta})$.
@@ 5.1 High-fidelity model simulation
Monitoring data: The mechanical model described in Section 2 represents an open drift located in the northern main drift of Gorleben for which monitoring data is available. The first convergence measurement took place on October 20, 1999, one day after the excavation of the drift. This is important because it allows for the measurement of primary creep, which is most pronounced at the beginning of the excavation. On the day of the initial measurement, the horizontal distance (1-3) measured $8.94\,m$, and the vertical distance (2-4) was $6.00\,m$, as depicted in Figure 2(a). The duration of the monitoring was approximately $14.3$ years, with a significant variation of time intervals between measurements. During the first two months, measurements were taken every two to three days in order to capture the high creep rates. Subsequently, the measurement interval was increased from one week to one month, and from the end of 2002 onwards, measurements were taken approximately every six months as a constant creep rate had been achieved. Since the measuring points of the convergence distances are anchored in the rock over an anchoring length, the deformations of the cavity are also measured at these fixed points in the rock salt. Those locations are defined as history points in the numerical model in FLAC3D, demonstrated in Figure 2(b). The convergences $\mathbf{y}_{\text{moni}}^{T}(t_{i})=\left[\Delta u_{x}(t_{i}),\,\Delta u_{z}(t_{i})\right]^{T}$ are measured as a vertical and horizontal distance between two points, therefore the displacements $u_{x}$ in horizontal and $u_{z}$ in vertical direction are obtained from the displacement vector $\mathbf{u}$ for the specific nodes. The uncertainty of the convergence measurements is around ±$\pm 0.5\,mm$, while observations are on the order $\mathcal{O}(10^{2})$. Consequently, we consider the monitoring data to be noise-free.
@@
Preliminary parameter selection: TUBSalt has a total of $25$ material parameters. Out of these, $5$ parameters can be read directly from laboratory stress-strain data, namely $K_{0}$, $G_{0}$, $\varepsilon_{v,d,b}$, $\sigma_{z,0}$ and $\rho$. In the following, these are considered as well-calibrated and independent of the monitoring data. The remaining $20$ parameters serve as potential parameters for calibration. We then conducted a preliminary, empirical study with only two simulations: several damage-associated strain components were deactivated, and the result was compared to a reference solution obtained with the full set of strain components (2.5). It has been found that the test case considered is dominated by primary and secondary creep mechanisms and damage-associated strain components are negligible. Details can be found in Appendix A. As a result, in the following we focus on a total of $7$ TUBSsalt parameters for primary and secondary creep: $\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$ and $E_{p}$ for primary creep and $\eta_{s}$, $\sigma_{0,eq,s}$ and $p_{s}$ for secondary creep, and thus $\bm{\theta}=\left[\eta_{p},\,\sigma_{0,eq,p},\,p_{p},\,E_{p},\,\eta_{s},\,\sigma_{0,eq,s},\,p_{s}\right]$. The parameter ranges of those $7$ parameters, as well as the parameters that are considered as fixed, are summarized in Appendix B, Table B1.
@@
Sampling of parameter values in Flac3D: Following the approach in Section 3.1, a parameter set for the constitutive model TUBSsalt based on laboratory strength and creep tests has been determined in a project presented in [50] for the Gorleben site. These values serve as characteristic parameter values for Gorleben salt and also help define reasonable parameter ranges for those parameters that need calibration, see also Table B1 in Appendix B. In order to create training and testing data by means of a multitude of simulations, the parameter values of the selected TUBSsalt parameters are sampled in the given ranges in Table B1 assuming a uniform distribution. For this purpose, a SciPy Quasi-Monte Carlo generator is used to generate a Sobol sequence by creating low-discrepancy, quasi-random numbers [65, 66]. The loop for parameter sampling is implemented in the creep analysis of Flac3D, so that the saved equilibrium state of the model is recalled for every iteration. The creep analysis is performed for the same duration of the monitoring. To maintain the state of equilibrium, the average ratio between out-of-balance and total forces is limited to $5\cdot 10^{-6}$. The timestep increases over time with a maximum of $\Delta t=5.56$ hours. With the given configuration, a single simulation run of the high-fidelity model requires approximately $4$ minutes.
@@
Processing the full-field simulation results to convergences: After carrying out the simulations to create training and testing data, the displacements derived at the history locations are converted into convergences at the time instances for which monitoring data are available. To determine the vertical convergences, the amounts of the vertical displacements $u_{z}(t_{i})$ at the top and bottom are summed up together. To determine the horizontal convergences, the amount of the horizontal displacements $u_{x}(t_{i})$ is multiplied by two to account for the symmetry of the system. The displacements of an exemplary full-field simulation at $t=14.5$ years is depicted in Figure 4. In order to compare the simulated convergences with the monitoring data, the convergences of the simulations must be set to zero at the time of the initial monitoring measurement. Therefore, the convergences from the first day after excavation of the drift can not be analysed and will be neglected. Convergences are determined at all points in time at which monitoring is available. As a result, Figure 5 shows the processed data of $200$ simulations. Figure 5a depicts the histograms obtained after sampling the input material parameters, while Figure 5b,c correspond to the vertical and horizontal convergences over time (blue lines) in comparison with the corresponding monitoring data (black squares) and the obtained probability density function (PDF) at selected time instances (red area). This figure illustrates how the uncertainty in material parameters propagates through the mechanical model implemented in FLAC3D, reflecting its nonlinearity. That is, uniformly sampled inputs result in heavily tailed outputs. The task of the GP-based surrogate is to represent this input-output relationship by learning the PDF of the convergences at different time instances as a function of the model parameters.
@@
Figure 5: Input-output simulation data at the monitoring location. a. Histograms for the input material parameters $\eta_{p}$, $E_{p}$, $\eta_{s}$, $p_{s}$, $\sigma_{0,eq,p}$, $p_{p}$, and $\sigma_{0,eq,s}$ of TUBSsalt. b. Vertical and c. horizontal convergence trajectories over time. b and c show the monitoring data (black squares) and $200$ model realizations (blue lines) from the FLAC3D simulations along with the obtained PDF at selected time instances (red area).